OpenStudy (anonymous):

Match the sequence (term) with the correct type of sequence (definition). Match Term Definition 128, 32, 8, 2, … A) Geometric, common ratio is 3 1, 3, 9, 27, … B) Geometric, common ratio is 0.25 5, 10, 15, 20, … C) Arithmetic, common difference is −3 20, 17, 14, 11, … D) Arithmetic, common difference is 5

OpenStudy (karolynxx):

128,32,8,2 .. 32/128 = 0.25 check.. 128 * 0.25 = 32 32 * 0.25 = 8 the common ratio is 0.25 (geometric) 1,3,9,27 3/1 = 3 check.. 1 * 3 = 3 3 * 3 = 9 9 * 3 = 27 the common ratio is 3 (geometric) 5,10,15,20 10 - 5 = 5 15 - 10 = 5 20 - 15 = 5 common difference is 5 (arithmetic) 20,17,14,11 20 - 3 = 17 17 - 3 = 14 14 - 3 = 11 common difference is -3 (arithmetic)